# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v3

@@ 1. Introduction
The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery or radiotherapy, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
@@
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019). Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, subsequently alleviating the burden on doctors and medical staff. To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
@@
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (RAG) over a Knowledge Base (KB) curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a KB update framework based on expert-provided edits to minimize future expert intervention. CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
@@
Our work aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use the terms ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators. (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous KB updates.
Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://github.com/microsoft/BYOeB. , designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs. (2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants at a tertiary eye hospital in India. Four doctors and two patient coordinators served as experts, actively verifying and correcting responses generated by the bot. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2.1. Patients’ Information Needs
Information seekers may also consult experienced patients within their social networks who can share valuable insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010). In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014). Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
@@ 2.2. Chatbots in Healthcare
Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023). They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023). Researchers have applied chatbots in various healthcare settings (Wutz et al., 2023), including mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023), appointment scheduling (Ittarat et al., 2023), seeking information for chronic disease management (Montagna et al., 2023), and assistance for new mothers (Yadav et al., 2019). Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk. Ada: https://ada.com/. Florence: https://florence.chat/., allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs. However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023). Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018). Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020), and few-shot learning. Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022). Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
@@
Finally, the use of AI in real-world healthcare deployments is limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and validated its utility to understand its integration into clinical workflows.
@@ 3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify key chatbot requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. From these interviews, we aimed to understand their current workflow, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps. Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience. They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. We conducted a thematic analysis to identify these key design requirements for the chatbot.
@@
Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance. To ensure accurate and contextual responses, our chatbot relies on a Knowledge Base (KB) curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
@@
Trustworthiness: Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or friends, they often confirm that information with their doctor. E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
@@
Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to busy schedules of doctors and coordinators, immediate responses are not always feasible. This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom KB curated by hospital staff. Subsequently, these responses gets reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. To ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
@@
Usability: After cataract surgery, patients are expected to reference their discharge summaries for any queries, but this can be challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role immediate family members play in monitoring the patient’s healthcare. Coupled with doctors’ input (“Hardly 10% of questions are asked by patients; attendants are the one who ask.”), this underscored the necessity for the bot to be usable by the four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators. To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal similar to (Jain et al., 2018; Findlater et al., 2009). We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022), despite the richer feature set of Telegram or the development flexibility of a custom chatbot. This choice aims to minimize the learning curve and ensure quick onboarding without requiring additional app installations. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s KB, minimizing similar edits in future.

### v4

@@ 1. Introduction
The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
@@
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019). Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, alleviating the burden on doctors and medical staff. To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
@@
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation over a knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a knowledge base update framework based on expert-provided edits to minimize future expert intervention. CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
@@
Our work aims to answer the following research questions: (RQ1) What is the role of CataractBot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and coordinators. (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous updates to the knowledge base. Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://github.com/microsoft/BYOeB. , designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs. (2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
@@ 2.1. Patients’ Information Needs
Information seekers may also consult experienced patients within their social networks who can share insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010). In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014). Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
@@ 2.2. Chatbots in Healthcare
Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023). They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023). Researchers have applied chatbots across diverse healthcare settings (Wutz et al., 2023), including scheduling appointments (Ittarat et al., 2023), collecting pre-consultation information (Li et al., 2024), assisting new mothers (Yadav et al., 2019), supporting individuals suffering from chronic diseases (Montagna et al., 2023; Hao et al., 2024), and offering mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023; Kim et al., 2024). Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk. Ada: https://ada.com/. Florence: https://florence.chat/., allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs. However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023). Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018). Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (Lewis et al., 2020), and few-shot learning. Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022). Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
@@
Finally, real-world healthcare deployments of AI are limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and understood its integration into clinical workflows.
@@ 3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. From these interviews, we aimed to understand current workflows, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps. Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience. They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. We conducted thematic analysis to identify key design requirements for the chatbot.
@@
Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance. To ensure accurate and contextual responses, our chatbot relies on a knowledge base curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
@@
Trustworthiness: Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or other sources, they often confirm that information with their doctor. E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
@@
Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to busy schedules of doctors and coordinators, immediate responses are not always feasible. This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom knowledge base curated by hospital staff. Subsequently, these responses get reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. o ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
@@
Usability: After cataract surgery, patients can reference their discharge summaries for any queries, but this is challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role family members play in monitoring the patient’s health. Coupled with doctors’ input (“Patients ask hardly 10% of the question… attendants are the ones who ask.”), this underscored the necessity for the bot to be usable by four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators. To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal, similar to prior work (Jain et al., 2018; Findlater et al., 2009). We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022) and therefore minimal learning curve and quick onboarding without additional installations, despite the richer feature set of other platforms like Telegram or the development flexibility of a custom chatbot. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s knowledge base, minimizing similar edits in future.